OpenStudy (anonymous):

I need to solve 2x^2 = 3x + 2. Can anyone show me how?

OpenStudy (anonymous):

using the quadratic formula

OpenStudy (anonymous):

2x^2 - 3x - 2 = 0

OpenStudy (anonymous):

just plug it in to the formula then?

OpenStudy (anonymous):

which part do you not understand?

OpenStudy (anonymous):

I guess I don't know the formula or understand how to use it

OpenStudy (anonymous):

for a quadratic equation in the form \[ax^2+bx+c\]x is \[(-b±sqrt(b^2-4ac))/2a\] You'll probably have two solutions, so use both the plus and the minus.

OpenStudy (anonymous):

okay \[x = \frac{-b \pm \sqrt{b^2 - 4*a*c}}{2a}\] thats the formula now you have 2x^2 - 3x - 2 = 0 ax^2 + bx + c = 0 plug in make sense?

OpenStudy (anonymous):

can I plug numbers into the equation?

OpenStudy (anonymous):

into the quad. formula

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

sure ^_^